A circle has its center at (2,-1) and one of a diameter at(5,1).Find the other end of the diameter.
Pls answer this..tnx.

Answer by stanbon(75887)

A circle has its center at (2,-1) and one end of a diameter is at(5,1).
Find the other end of the diameter.
-------
Let the other end be (x,y)
---
Solve for x and for y:
(x+5)/2 = 2
x+5 = 4
x = -1
----------------
(y+1)/2 = -1
y+1 = -2
y = -3
--------
Other end: (-1,-3)
====================
cheers,
Stan H.
=====================